# Predicting and Understanding College Student Mental Health with Interpretable Machine Learning

### v2

@@ 5.3. Scoring for Interaction Labels
To analyze how each feature contributes to scoring for interaction labels, we visualize the importance of relevant features across all 217 students. Figure 8 displays the top features (marked with a star) for each interaction label:
@@
As illustrated in Algorithm 2, the scores for each interaction label are calculated based on the relative importance of various features, measured against a threshold value derived from the mean of distribution across all individuals. Next, we show the thresholds for Leisure Time and Sleep to demonstrate how feature contributions influence the final label scoring.
@@
In Figure 9, the dotted lines highlight the threshold values for the most important features (F4, F13, F3, F5, F14, and F8) for Leisure Time . F4 reflects walking duration: if a student’s walking duration is less than the threshold (3 hours), it indicates leisure activity. Consequently, the LeisureScore is incremented proportionally to the feature importance of walking for the respective student. Similarly, F8 considers voice detection in others’ dormitories. If the duration exceeds the threshold (15 minutes), it suggests that the student is having leisure time with friends, causing an increase in LeisureScore by the value of importance of the corresponding feature.
In Figure 10, the dotted lines highlight the threshold values for the most important features (F6, F4, F7, and F1) for Sleep. F6 indicates total sleep duration: if a student’s sleep exceeds 7 hours, representing the standard benchmark for healthy sleep, the SleepScore is incremented based on the student’s feature importance for sleep. F4, which measures the time spent in the room, shows that durations greater than 12.9 hours typically signify overnight dorm occupancy, often indicative of sleep rather than other activities. F7 corroborates healthy sleep with thresholds above 7.5 hours, while F1 reflects quietness, with values below 6 minutes suggesting an environment conducive to rest.
@@ 6. Related Work
7. Conclusion
This paper addresses the challenges of predicting and understanding mental health among college students. Using a cutting-edge longitudinal dataset focused on college student behavior, we present I-HOPE, the first hierarchical model designed to predict and interpret mental health status at the individual level. Our evaluation shows that I-HOPE significantly outperforms baseline methods, offering valuable insights into how various behavioral features influence mental health outcomes. This research lays the foundation for future efforts aimed at predicting and understanding the mental health of college students. We hope that our findings inspire further exploration and innovation in this critical area of study, ultimately contributing to enhanced support systems for student wellness.

### v3

@@ 5.3. Scoring for Interaction Labels
To analyze how each feature contributes to scoring for each interaction label, we visualize the feature importances of relevant features derived from random forest models across all individuals. Figure 8 displays the top features (marked with a star) for each interaction label:
@@
As illustrated in Algorithm 2, the scores for each interaction label are calculated based on the weighted importance of various features (i.e., NWFI in Algorithm 2), measured against a threshold value derived from the mean of distribution across all individuals. Next, we show the thresholds for Leisure Time and Sleep to demonstrate how feature contributions influence the final label scoring.
@@
In Figure 9, the dotted lines highlight the threshold values for the most important features (F4, F13, F3, F5, F14, and F8) for Leisure Time . F4 reflects walking duration: if a student’s walking duration is less than the threshold (3 hours), it indicates leisure activity. Consequently, the LeisureScore is incremented proportionally to the weighted feature importance of walking for the respective student. Similarly, F8 considers voice detection in others’ dormitories. If the duration exceeds the threshold (15 minutes), it suggests that the student is having leisure time with friends, causing an increase in LeisureScore by the corresponding weighted feature importance.
In Figure 10, the dotted lines highlight the threshold values for the most important features (F6, F4, F7, and F1) for Sleep. F6 indicates total sleep duration: if a student’s sleep exceeds 7 hours, representing the standard benchmark for healthy sleep, the SleepScore is incremented based on the student’s weighted feature importance for sleep. F4, which measures the time spent in the room, shows that durations greater than 12.9 hours typically signify overnight dorm occupancy, often indicative of sleep rather than other activities. F7 corroborates healthy sleep with thresholds above 7.5 hours, while F1 reflects quietness, with values below 6 minutes suggesting an environment conducive to rest.
@@ 6. Related Work
7. Conclusion and Future Work
This paper addresses the challenge of personalized predictions and understanding mental health among college students. Leveraging a cutting-edge longitudinal dataset, we present I-HOPE, the first hierarchical model designed to predict and interpret mental health status at the individual level. I-HOPE bridges the gap between black-box machine learning and actionable insights, offering both accurate and interpretable, personalized predictions. These contributions are especially important in the mental health domain, where individualization is critical.
Our evaluation shows that I-HOPE has a better performance than baseline methods and offers insights into how behavioral patterns, such as sleep, phone usage, and social time, relate to mental health outcomes on a personalized level. However, like any data-driven model, there are several areas that offer opportunities for further improvement, which we describe as follows.
From a data perspective, our analysis relied on features aggregated at a daily level. Incorporating higher-resolution hourly data could reveal more granular behavioral patterns, potentially enhancing prediction accuracy. Additionally, certain data processing decisions—such as replacing missing values with zeros—may oversimplify real-world behaviors. Future work could explore more sophisticated imputation techniques to better capture underlying trends.
From a modeling perspective, while the current framework uses five interaction labels—empirically motivated and preliminarily validated—future work could systematically explore how varying the number of labels impacts both predictive performance and interpretability. Additionally, incorporating aggregate metrics (e.g., mean performance) and confidence intervals during evaluation would offer a more robust assessment of model reliability.
Furthermore, the lack of a fixed random seed introduced unintended variability in reported results; standardizing this in future experiments will improve reproducibility and consistency.
These refinements present a pathway for advancing mental health research, enabling the development of more interpretable, personalized, and actionable models.